OpenStudy (gottennis121):

Medal! (Pythagorean Theorem) The legs of a right triangle measure 24 feet and 45 feet. What is the length of the hypotenuse? A. 51 feet B. 69 feet C. 38.07 feet D. 8.31 feet. Thank you! @josedavid @sleepyjess @pooja195 @undeadknight26

pooja195 (pooja195):

A^2+B^2=C^2

pooja195 (pooja195):

There should be a diagram with this =/

OpenStudy (gottennis121):

Yea, I know the formula, but got the answer 2,601.

pooja195 (pooja195):

@Nnesha

Nnesha (nnesha):

hypotenuse is c which is big side of triangle

Nnesha (nnesha):

yes you got 2601 then you have to take square root bec \[24^2 + 45^2 = C^2\] \[c^2 =2601\]

pooja195 (pooja195):

oh so then get the square root of 2601

OpenStudy (gottennis121):

I got it. \[24 x 24 + 45 x 45 = \sqrt2601 = 51.

Nnesha (nnesha):

yep!!

Nnesha (nnesha):

gO_Od job!

OpenStudy (gottennis121):

:D

Nnesha (nnesha):

D: